OpenStudy (anonymous):

differentiate 1/sqrtx

OpenStudy (anonymous):

good problem, how can you re-write this to make it seem easy to compute?

OpenStudy (anonymous):

can you perhaps write this as an expoenent

OpenStudy (anonymous):

the denominator

OpenStudy (anonymous):

\[\frac{1}{\sqrt{x}}=x^{-\frac{1}{2}}\] is what they were trying to get you to write. now use power rule

OpenStudy (anonymous):